OpenStudy (anonymous):

Find the axis of symmetry of the parabola formed by the following quadratic function: y = -5x2 + 7x - 9 A. x = -7/5 B. x = 7/5 C. x = -7/10 D. x = 7/10

OpenStudy (anonymous):

its D

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

for the parabola y = ax^2+bx + c, axis of symmetry is x = -b/2a

OpenStudy (anonymous):

i knew it was something like that.! i left my notes at a friends house & started a test online & i know very few of the questions :(